AshtonSpaysky:

In right traingle LMN, L and M are complementary angles and sin(L) is . What is cos(M)?

TheSmartOne:

You need to know this: sin (90-x) = cos (x) Since L and M are complementary angles, that means sin(L) = cos(M) Hopefully you can now find your answer (: